OpenStudy (anonymous):

Which of these ordered pairs is a solution to the linear inequality y – 4x ≤ –6? A. (–2, 4) B. (1, –2) C. (1, 3) D. (2, 3)

OpenStudy (mathstudent55):

Plug in each ordered pair until you find one that makes the inequality true.

OpenStudy (anonymous):

B (-2) - (4*(1)) = -6 -6 = -6